OpenStudy (anonymous):

simplify square root of 2 (square root of 6)

OpenStudy (anonymous):

sqrt2 * sqrt6 = sqrt12

OpenStudy (anonymous):

can that me simplified

OpenStudy (anonymous):

be not me

OpenStudy (anonymous):

yes 2sqrt3

OpenStudy (anonymous):

yes it can nvm

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it can 6 times the square root 2? (6^1/2) * ( 2^1/2) = (6*2)^1/2 = 12^1/2 = (3 *4) ^1/2 = 2(3^1/2)

OpenStudy (anonymous):

ok thank you